A person in a kayak starts paddling, and it accelerates from 0 to 0.798 m/s in a distance of 0.526 m. If the combined mass of the person and the kayak is 63.8 kg, what is the magnitude of the net force acting on the kayak?
Answer #1

From kinematic equation

v^2=vo^2+2*a*x

a =[(0.798 m/s)^2 -(0 m/s)^2]/(2*0.526 m)

a =0.605326996 m/s^2

From Newton's 2nd law

Fnet =m*a

Fnet =63.8 kg *0.605326996 m/s^2 =38.61986236 N

Fnet =38.6 N